OpenStudy (anonymous):

You are hired to help minimize production costs. Your job is to choose the form of packaging for a new product from the five choices below. The package must hold at least 1000 cubic inches. The cost of the cardboard for the packaging is $0.03 per square inch. Your project should include the volume, total surface area, and materials cost for each solid given below, including the formulas you used and each step of your work. Make sure to use the formulas given in your lessons and round your answers to the nearest hundredth.

OpenStudy (anonymous):

Also, this goes along with it. :) Solid 1: Square Prism with each side of the base equal to 10 in. and a height of 8 in. Solid 2: Square Pyramid with each side of the base equal to 13 in. and a height of 21 in. Solid 3: Cylinder with a radius of 7 in. and a height of 8 in. Solid 4: Cone with a radius of 10 in. and a height of 9 in. Solid 5: Sphere with a radius of 7 in.

OpenStudy (janu16):

didnt you just post this like 30 minutes ago?

OpenStudy (anonymous):

Yes, but nobody would answer me. @Janu16

OpenStudy (janu16):

@rebeccaxhawaii

OpenStudy (janu16):

@jhonyy9

OpenStudy (michele_laino):

hint: we have to choose the solid shape which has the least total surface

OpenStudy (anonymous):

Okay. I have my formulas, but I need to understand how to plug in the info into the formula. Could you help me with that?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Thank you! :)

OpenStudy (michele_laino):

for example, if we take the sphere with a radius of 7 meters, we have this volume: \[V = \frac{{4\pi }}{3}{r^3} = \frac{{4 \times 3.14}}{3} \times {7^3} = 1436{m^3}\]

OpenStudy (michele_laino):

and the total Surface, is: \[S = 4\pi {r^2} = 4 \times 3.14 \times {7^2} = 615.44{m^2}\]

OpenStudy (anonymous):

That's easy to understand so far.

OpenStudy (michele_laino):

so, please take a note of that surface

OpenStudy (anonymous):

Okay.

OpenStudy (michele_laino):

next let's consider the solid 4), namely the cone

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

When I do these, am I just doing the total surface area or do I need to include the lateral as well?

OpenStudy (michele_laino):

here the volume is: \[{V_4} = \frac{{\pi {r^2}h}}{3} = \frac{{3.14 \times {{10}^2} \times 9}}{3} = ...inche{s^3}\]

OpenStudy (michele_laino):

I think that we have to compute the total Surface, since we have to mi nimize the cost of cardboard for packaging

OpenStudy (anonymous):

Okay

OpenStudy (michele_laino):

please complete

OpenStudy (michele_laino):

what is \(V_4\)?

OpenStudy (anonymous):

210.6?

OpenStudy (anonymous):

or would it be 842.4?

OpenStudy (michele_laino):

I got \(V_4=942\)

OpenStudy (anonymous):

Okay, that's probably the correct answer.. I didn't round like it stated, I'm sorry lol

OpenStudy (michele_laino):

ok! Next we have to compute the total Surface of such cone

OpenStudy (anonymous):

Okay! :) also, thank you so much for helping. This means so much!

OpenStudy (michele_laino):

the total Surface \(S_4\) of such cone, is: \[\begin{gathered} {S_4} = \pi {r^2} + \frac{{2\pi r\sqrt {{r^2} + {h^2}} }}{2} = \hfill \\ \hfill \\ = 3.14 \times {10^2} + \frac{{2 \times 3.14 \times 10 \times \sqrt {{{10}^2} + {9^2}} }}{2} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

This is probably wrong, but I got 3171.4... Ha....

OpenStudy (michele_laino):

more explanation: